# Training-Free Bias Mitigation by LLM-Assisted Bias Detection and Latent Variable Guidance

### v3

@@ 6.1 Experimental Settings
We employ the Parti Prompt dataset (yu2022scaling), which comprises over 1,600 diverse English prompts intended to rigorously evaluate T2I generation models and probe their limitations. In this work, we focus on the uniform prompt distribution, as no particular target distribution is assumed. For practical efficiency, we randomly sample 50 prompts from the full dataset as in section 4. The selected subset of prompts is listed in subsection B.1.
@@
We compare FairT2I against classifier-free guidance (CFG) (ho2022classifier) at three guidance scales (7.0, 4.0, 1.0), which modulate the trade‑off between image fidelity and diversity.
@@
We use Stable Diffusion 3.5-large (sd3) as the T2I model. We employ Claude 3.7-sonnet (anthropic2025claude37sonnet) for LLM-assisted bias detection, and GPT-4o-mini (gpt4) to fuse textual input $\mathbf{y}$ with sensitive attribute $\mathbf{z}$. Generation hyperparameters and LLM prompts are in subsection D.2.
@@ 6.2 Automatic Evaluation
To assess image fidelity, we compute the Fréchet Inception Distance (FID) (heusel2017gans) on the COCO Karpathy test split (lin2014microsoft). To evaluate text–image alignment, we report CLIPScore (hessel2021clipscore) between each generated image and its input prompt. To quantify diversity, we embed all generated images into CLIP (clip) and BLIP2 (li2023blip) feature spaces, compute the covariance of embeddings, and use its trace as the diversity metric. For each prompt, we generate 200 images and compute these metrics accordingly.
@@ 8 Limitations
FairT2I reflects latent attributes in image generation via explicit prompt modifications. As demonstrated by hirota2024saner, information such as gender that is explicitly specified is accurately captured by the text encoder in the generated images. However, long prompts and complex logical relations (tang2023lemons; zhang2023iti) may not be reflected in the outputs. This limitation could potentially be addressed by using a long-context model (zhang2024long) or a vision-language model-based text encoder (chen2025multimodal).
@@ B.1 Evaluation Dataset Details
For efficient evaluation, we used a subset of 50 randomly selected prompts from each dataset: profession names from the Stable Bias dataset [Luccioni_2023] and textual prompts from the Parti Prompt dataset [yu2022scaling]. The selected subsets are listed in Table 9 and Table 11, respectively.
@@ B.2 Pilot Study Details
We conducted a pilot comparison of four LLMs: GPT-4o gpt4, Claude-3.7-Sonnet anthropic2025claude37sonnet, LLaMA-3.3 grattafiori2024llama3herdmodels, and DeepSeek-V3 liu2024deepseek, to assess their ability to detect potential bias-relevant categories and attributes in text prompts. The comparison was performed using two evaluation subsets: 50 prompts sampled from the Stable Bias dataset and another 50 prompts from the Parti Prompt dataset, as listed in subsection B.1.
@@ B.3 User Study Details
We prepared a total of 100 text prompts, consisting of 50 prompts from the Stable Bias dataset [Luccioni_2023] and 50 prompts from the PartiPrompts dataset [yu2022scaling], as detailed in subsection B.1. To maintain reasonable annotation time, these 100 prompts were divided into five separate tasks, each containing 20 prompts. Different tasks were assigned to different participants to avoid repetition.
@@ C.3 Fair Diffusion Setup Details
Except for these settings, all other parameters matched those of the original work friedrich2023fair.
@@ D.1 Evaluation Metrics Details
We calculated FID heusel2017gans to assess image fidelity against the COCO Karpathy test split [lin2014microsoft]. We used the PyTorch implementation of FID from the pytorch-fid library [^7] https://github.com/mseitzer/pytorch-fid.
@@
To evaluate text-image alignment, we computed CLIPScore hessel2021clipscore.
@@
To quantify diversity using CLIP clip features, we calculated the trace of the covariance matrix of image embeddings.
@@
Similar to CLIP Trace, we also quantified diversity using BLIP-2 li2023blip features.
@@ D.5 User Study Details
We sampled 50 prompts from the Parti Prompts dataset [yu2022scaling], as detailed in subsection B.1. To keep the annotation workload manageable, these 50 prompts were divided into five separate tasks, each containing 10 prompts. Each task was assigned to a distinct group of participants to prevent repetition across annotators.

### v4

@@ 6.1 Experimental Settings
We employ the Parti Prompt dataset (Yu et al., 2022), which comprises over 1,600 diverse English prompts intended to rigorously evaluate T2I generation models and probe their limitations. In this work, we focus on the uniform prompt distribution, as no particular target distribution is assumed. For practical efficiency, we randomly sample 50 prompts from the full dataset as in section 4. The selected subset of prompts is listed in subsection B.1.
@@
We compare FairT2I against classifier-free guidance (CFG) (Ho and Salimans, 2022) at three guidance scales (7.0, 4.0, 1.0), which modulate the trade‑off between image fidelity and diversity.
@@
We use Stable Diffusion 3.5-large (Esser et al., 2024) as the T2I model. We employ Claude 3.7-sonnet (Anthropic, 2025) for LLM-assisted bias detection, and GPT-4o-mini (OpenAI et al., 2024) to fuse textual input $\mathbf{y}$ with sensitive attribute $\mathbf{z}$. Generation hyperparameters and LLM prompts are in subsection D.2.
@@ 6.2 Automatic Evaluation
To assess image fidelity, we compute the Fréchet Inception Distance (FID) (Heusel et al., 2017) on the COCO Karpathy test split (Lin et al., 2014). To evaluate text–image alignment, we report CLIPScore (Hessel et al., 2021) between each generated image and its input prompt. To quantify diversity, we embed all generated images into CLIP (Radford et al., 2021) and BLIP2 (Li et al., 2023) feature spaces, compute the covariance of embeddings, and use its trace as the diversity metric. For each prompt, we generate 200 images and compute these metrics accordingly.
@@ 8 Limitations
FairT2I reflects latent attributes in image generation via explicit prompt modifications. As demonstrated by Hirota et al. (2024), information such as gender that is explicitly specified is accurately captured by the text encoder in the generated images. However, long prompts and complex logical relations (Tang et al., 2023; Zhang et al., 2023) may not be reflected in the outputs. This limitation could potentially be addressed by using a long-context model (Zhang et al., 2024a) or a vision-language model-based text encoder (Chen et al., 2025b).
@@ B.1 Evaluation Dataset Details
For efficient evaluation, we used a subset of 50 randomly selected prompts from each dataset: profession names from the Stable Bias dataset [Luccioni et al., 2023] and textual prompts from the Parti Prompt dataset [Yu et al., 2022]. The selected subsets are listed in Table 9 and Table 11, respectively.
@@ B.2 Pilot Study Details
We conducted a pilot comparison of four LLMs: GPT-4o OpenAI et al. [2024], Claude-3.7-Sonnet Anthropic [2025], LLaMA-3.3 Grattafiori et al. [2024], and DeepSeek-V3 Liu et al. [2024], to assess their ability to detect potential bias-relevant categories and attributes in text prompts. The comparison was performed using two evaluation subsets: 50 prompts sampled from the Stable Bias dataset and another 50 prompts from the Parti Prompt dataset, as listed in subsection B.1.
@@ B.3 User Study Details
We prepared a total of 100 text prompts, consisting of 50 prompts from the Stable Bias dataset [Luccioni et al., 2023] and 50 prompts from the PartiPrompts dataset [Yu et al., 2022], as detailed in subsection B.1. To maintain reasonable annotation time, these 100 prompts were divided into five separate tasks, each containing 20 prompts. Different tasks were assigned to different participants to avoid repetition.
@@ C.3 Fair Diffusion Setup Details
Except for these settings, all other parameters matched those of the original work Friedrich et al. [2023].
@@ D.1 Evaluation Metrics Details
We calculated FID Heusel et al. [2017] to assess image fidelity against the COCO Karpathy test split [Lin et al., 2014]. We used the PyTorch implementation of FID from the pytorch-fid library [^7] https://github.com/mseitzer/pytorch-fid.
@@
To evaluate text-image alignment, we computed CLIPScore Hessel et al. [2021].
@@
To quantify diversity using CLIP Radford et al. [2021] features, we calculated the trace of the covariance matrix of image embeddings.
@@
Similar to CLIP Trace, we also quantified diversity using BLIP-2 Li et al. [2023] features.
@@ D.5 User Study Details
We sampled 50 prompts from the Parti Prompts dataset [Yu et al., 2022], as detailed in subsection B.1. To keep the annotation workload manageable, these 50 prompts were divided into five separate tasks, each containing 10 prompts. Each task was assigned to a distinct group of participants to prevent repetition across annotators.